# A Random Rule Model

### 04 Mar 2026 v1

@@ 8 Mechanisms
We now examine the behavioral mechanisms underlying rule-gating’s predictive performance. The responsibility weights $w_{f}$ and context-dependent patterns reported below are population-level objects: they describe how the fitted mixture varies across menus in a dataset aggregating choices across many individuals. This is consistent with within-individual variation in rule use, between-individual heterogeneity, and noise. Without individual-level panel data, these components cannot be separated; mechanism-level statements should be read as describing population-level reallocation.
8.1 Responsibility weights $w_{f}$
Two patterns stand out. First, the attention rules A1 and A2 absorb the two largest shares of effective responsibility ($w_{\textsc{A1}}\approx 0.193$ and $w_{\textsc{A2}}\approx 0.146$), with a combined attention weight of about $34\%$. This is expected: the attention rules are decisive on essentially all menus and serve as a catch-all channel. Among the other rules, MMn (security, $15.7\%$), MMx (aspiration, $11.4\%$), SAL (most salient comparison, $10.7\%$), and MAP (modal payoff, $8.5\%$) carry the largest responsibility weights. Second, because multiple rules are typically active on the same menu (on average, about $5.8$ rules are both active and consistent with the majority choice direction), these responsibility weights reflect the outcome of competition among substitutable procedures rather than mechanical assignment to the only available rule.
It is useful to separate the library into two conceptual channels. The structured rules (MMn, MMa, MMx, MAP, SAL, SAL2, REG, REGmed, DIS, DISmed) encode distinct behavioral mechanisms that generate non-trivial, menu-dependent recommendations. The attention channel (A1 and A2) captures inattention, hasty responding, or trembling: A1 always recommends left and A2 always recommends right, regardless of lottery content. These attention rules absorb about one-third of effective responsibility, reflecting the prevalence of default-like behavior in the online (MTurk) environment. Because the marginal left-choice frequency is $\bar{p}\approx 0.510$ (essentially one-half), this weight cannot be explained by aggregate position bias; rather, the gate allocates attention weight in a menu-dependent way. Part of this weight may, however, reflect absorption of residual prediction error by always-active rules rather than a cleanly identified behavioral channel. Our interpretability claims concern the structured rules conditional on the inattention channel. Section 9.4 confirms that mechanism patterns persist when the attention channel is removed entirely.
To summarize how concentrated effective responsibility is, we report
$\mathrm{HHI}(F)=\sum_{f\in\mathcal{F}}w_{f}^{2},\qquad N_{\mathrm{eff}}(F)=\frac{1}{\mathrm{HHI}(F)}.$
In our estimates, $\mathrm{HHI}(F)\approx 0.122$ and $N_{\mathrm{eff}}(F)\approx 8.2$ (see Figure 4). For comparison, uniform weights across all $12$ rules would give $\mathrm{HHI}=1/12\approx 0.083$ and $N_{\mathrm{eff}}=12$. The observed concentration is therefore moderately above the uniform benchmark — driven primarily by the attention channel — but far from the winner-take-all extreme ($\mathrm{HHI}=1$).
8.2 Rule concentration
(a) Predictive ablation index $\phi(f)$.
(b) Concentration Impact $\sigma_{N}(f)$.
We next quantify which procedures are necessary for predictive fit. Figure 5(a) reports the ablation index $\phi(f)$, the relative change in test MSE when rule $f$ is removed from the full library and the gate is retrained on the reduced set. We ablate each of the 10 behavioral rules (excluding attention rules A1/A2, which serve a structural role ensuring decisiveness at every menu). A small subset of rules drives most of the predictive gains: removing SAL2 or SAL produces the largest deterioration, followed by MMx, MMn, and REGmed. Notably, SAL2 has a relatively high ablation impact despite carrying a lower average responsibility weight than MMn or SAL; this illustrates the distinction between $w_{f}$ (average usage) and $\phi(f)$ (marginal predictive contribution), and suggests that the second-most-salient comparison provides incremental content not replicated by other rules. By contrast, removing MAP, MMa, REG, or DIS/DISmed has little effect on MSE, suggesting that these rules are largely redundant given the rest of the library (or apply primarily in regions of menu space that contribute little to aggregate loss).
Ablations identify rules that matter for accuracy, but they do not indicate whether a rule acts as a unifier (concentrating effective responsibility on a small repertoire) or as a substitute that diversifies the fitted mixture. To capture this, in Figure 5(b), we report the relative change in the effective number of rules $\sigma_{N}(f)$. We find that $\sigma_{N}(f)$ is negative for all non-attention rules: when any single rule is removed, the refitted gate reallocates responsibility more sharply onto a smaller set of remaining procedures. The magnitude is largest for the same procedures that have large ablation impacts, indicating that these rules both attract mass and shape the overall mixture structure.
Overall, predictive accuracy is driven by a small set of procedures—salience-based comparisons and extreme-outcome comparisons—as evidenced by their large $\phi(f)$, while the responsibility weights $w_{f}$ additionally highlight modal-outcome comparisons (MAP) as receiving substantial mass despite a modest ablation impact. The remaining rules do not simply “sit idle”: their negative $\sigma_{N}(f)$ values show that they act as substitutes that diversify the fitted mixture. Appendix Table A.4 reports split-variability intervals for $\phi(f)$, confirming that the main rankings are stable. Such a decomposition is not available in the more flexible predictors that outperform rule-gating, and it provides a transparent characterization of which bounded-rationality mechanisms are empirically relevant in a large dataset of choice under risk.
8.3 Drivers of rule selection
Complexity
and the associated excess dissimilarity
$d(A)\;:=\;\Delta_{\mathrm{CDF}}(A)-\big|EV(L^{1})-EV(L^{2})\big|.$
Figure 6 shows clear and interpretable shifts in rule selection along tradeoff complexity. As $\mathrm{TC}(A)$ increases, the two attention rules (A1/A2) display a pronounced increase in mean effective weight, consistent with the view that more complex menus are associated with a greater reliance on coarse inattention in the learned CA-RRM. Among the structured procedures, MMx (aspiration/extreme-max comparison) rises with complexity, whereas MAP (modal payoff comparison) declines; SAL and MMn are comparatively stable across deciles. Overall, increasing tradeoff complexity appears to tilt the fitted mixture away from modal-outcome comparisons and toward more extreme-outcome and attention-based procedures.
The same exercise for the full library (Appendix Figure A.3) reveals additional structure among secondary rules: in particular, regret- and disappointment-type procedures (REG/REGmed and DIS/DISmed) exhibit a decline in weight as complexity increases, suggesting that these more evaluative comparative procedures play a smaller role in high-tradeoff menus relative to simpler screening and extreme-outcome mechanisms.
Risk
A second natural driver is the risk profile of the menu. Different procedures emphasize different aspects of risk, such as dispersion, downside exposure, and sensitivity to rare events. As a first-pass risk descriptor available directly from the menu, we consider the signed asymmetry in dispersion across the two lotteries,
$\mathrm{RiskAsym}(A)\;:=\;\sigma(L^{1})-\sigma(L^{2}),$
where $\sigma(L)$ denotes the standard deviation of lottery $L$. This is a directional measure: positive values indicate that $L^{1}$ is more dispersed, and negative values indicate that $L^{2}$ is. Figure 7 plots mean responsibility weights across deciles of this signed difference, so that the $x$-axis runs from menus where $L^{2}$ is riskier (left) to menus where $L^{1}$ is riskier (right). We use the absolute value $|\mathrm{RiskAsym}(A)|$ when referring to the magnitude of dispersion asymmetry without regard to direction.
Figure 7 provides a complementary view based on dispersion asymmetry. Along deciles of $\mathrm{RiskAsym}(A)$, the fitted mixture shifts away from extreme-outcome comparisons: both MMn (security) and MMx (aspiration) show a clear decrease in mean weight as risk asymmetry increases. By contrast, the attention rules and SAL remain relatively flat, while MAP displays an increasing pattern. Thus, in menus where dispersion is more asymmetric across the two options, the gate reallocates mass away from max/min-type screening and toward modal-outcome comparisons, with attention and salience mechanisms playing a more stable background role. More secondary rules show less systematic movements with risk (Appendix Figure A.4).
These diagnostics show that rule selection varies systematically with salient menu primitives: along complexity, the mixture tilts toward inattention and extreme-outcome screening and away from modal-outcome comparisons; along dispersion asymmetry, the shifts reverse on key dimensions. Menu-dependent selection across a small repertoire of procedures is economically meaningful.
9 Robustness
The RRM framework makes several modeling choices that could, in principle, drive its empirical performance: the composition of the rule library, the FOSD-based activation criterion, the use of interpretable gate features, and the softmax gating architecture itself. This section subjects each of these choices to targeted robustness checks. We begin with the identification diagnostics and decomposition stability (Section 9.1), then document library coverage, before turning to directed tests: cross-fitted library selection, attention-rule removal, variation of the activity discipline, and three further checks (placebo test, feature-based baselines, raw-encoding gate) reported in the appendix.
@@ 9.1 Identification diagnostics
$9,823$ of $9,831$ menus (99.999.9%) are two-sided, confirming that the log-odds representation (9) is well-defined on virtually the entire sample.
@@
Table 4 reports the empirical support for conditions (G1) and (G2) of Theorem 1, evaluated on the baseline gate features $z(A)\in\mathbb{R}^{12}$. In the Choices13k data, the $12$-dimensional features are essentially unique across menus, so exact-equality cells are singletons. We supplement exact equality with $k$-means clustering on $z(A)$, grouping menus into $50$ cells. [^10] The number of clusters ($50$) is chosen so that each cell contains at least $|\mathcal{F}|-1$ menus while keeping within-cell feature variation modest. The result is qualitatively stable to alternative cluster counts (e.g., $30$, $75$, $100$). Of these, $41$ cells achieve $\mathrm{rank}(H^{(k)})=|\mathcal{F}|-1$, well above the $d_{\mathrm{eff}}+1=12$ needed for (G1). Among the G1-passing cells, the augmented centroid matrix has full affine rank, satisfying (G2). Both conditions are supported. The rank conditions are structurally stable: because the relevant singular values are well above the SVD tolerance ($10^{-8}\times\sigma_{1}$), small perturbations in $\hat{r}(A)$ do not change the ranks.
Table 4: Global identification verification (D3). Conditions (G1) and (G2) of Theorem 1 are checked on the Choices13k data using $k$-means clustering on interpretable $z$-features ($d=11$).
@@
Fold-level stability of the decomposition objects is reported in the appendix (Table A.5): the rule rankings by $w_{f}$ and $\phi(f)$ are highly stable across independent training samples.
9.2 Library Coverage
@@
It is useful to document that predictive fit is not driven by a knife-edge mapping from menus to a single decisive procedure. For each menu, let $N_{\mathrm{cons}}(A)$ count the number of rules that are simultaneously active and consistent with the observed majority choice direction. Figure 8 reports the distribution across the 9,8319{,}831 menus. Menus that pin down a unique decisive rule are rare ($\Pr[N_{\mathrm{cons}}(A)=1]\approx 0.16\%$), while overlap is substantial on average ($\mathbb{E}[N_{\mathrm{cons}}(A)]\approx 5.8$). Moreover, the average number of active rules per menu is about $9.5$ out of $12$, so activity rarely collapses the library to a small subset. These diagnostics confirm that the gate is learning selection among genuinely competing procedures rather than exploiting mechanically unique recommendations.

### 19 Mar 2026 v2

@@ 7 Application: Procedural Heterogeneity in Risky Choice
This section estimates procedural heterogeneity in risky choice using the choices13k dataset. We first verify the identification conditions on the data and establish out-of-sample predictive performance. We then present the estimated behavioral decomposition - responsibility weights, ablation indices, and concentration measures - followed by an econometric cross-check via the two-step estimator. Finally, we report comparative statics in rule responsibility along menu primitives and portability to an independent dataset. The responsibility weights $w_{f}$ and context-dependent patterns reported below are population-level objects: they describe how the fitted mixture varies across menus in a dataset aggregating choices across many individuals. Without individual-level panel data, within-individual and between-individual components cannot be separated; mechanism-level statements should be read as describing population-level reallocation.
@@ 7.1 Identification diagnostics
$9,831$ of $9,831$ menus (100.0100.0%) are two-sided, confirming that the log-odds representation (9) is well-defined on virtually the entire sample.
@@
Table 2 evaluates the rank conditions (G1) and (G2) of Corollary 2 on the baseline gate features $z(A)\in\mathbb{R}^{12}$. In the Choices13k data, the $12$-dimensional features are essentially unique across menus, so exact-equality cells are singletons. We supplement exact equality with $k$-means clustering on $z(A)$, grouping menus into $50$ cells. [^8] The number of clusters ($50$) is chosen so that each cell contains at least $|\mathcal{F}|-1$ menus while keeping within-cell feature variation modest. Of these, $41$ cells achieve $\mathrm{rank}(H^{(k)})=|\mathcal{F}|-1$, well above the $d_{\mathrm{eff}}+1=12$ needed for (G1). Among the G1-passing cells, the augmented centroid matrix has full affine rank, satisfying (G2). Both conditions are supported under approximate binning. The rank conditions are structurally stable: because the relevant singular values are well above the SVD tolerance ($10^{-8}\times\sigma_{1}$), small perturbations in $\hat{r}(A)$ do not change the ranks.
We note that the $k$-means diagnostic rests on an approximation that lies outside the formal identification theory: Theorem 1 requires menus in a cell to share the exact feature value $x$, whereas clustering only ensures approximate homogeneity within each cell. This is a standard finite-sample device analogous to binning in semiparametric estimation, and its validity depends on $\omega(x)$ varying smoothly over the feature support. The overidentification $J$-test reported in Section 7.8 provides a complementary, estimation-based check on the adequacy of the gate restriction that does not rely on exact-equality cells.
Table 2: Global identification verification (D3). Conditions (G1) and (G2) of Corollary 2 are checked on the Choices13k data using $k$-means clustering on interpretable $z$-features ($d=11$).
@@
7.2 Predictive validation
Before examining the behavioral decomposition, we verify that the MSE-fitted model predicts well out of sample. Table 3 reports mean test MSE across 50 random train/test splits, with learning-rate selection as described in Section 6. The rule-gating predictor substantially outperforms the standard structured models (Neural EU/PT/CPT and Value-Based) and the flexible Context-Dependent model, closing nearly all of the gap to the strongest benchmark (MOT), providing strong support for the identified decomposition.
Concretely, rule-gating improves on the Value-Based model and on all Neural EU/PT/CPT baselines by a substantial margin. These gains indicate that a small repertoire of transparent “lines of conduct,” combined with a flexible gate that learns when each procedure applies, captures substantial regularity in the data that is missed by standard preference-based architectures. Rule-gating also substantially outperforms the flexible Context-Dependent model. Only the Mixture of Theories (MOT) achieves a marginally lower MSE, with the gap closing to about $2.5\%$ of the MSE level; this is consistent with the fact that MOT allows richer menu interactions than any fixed rule library, but indicates that the structured decomposition captures nearly all of the exploitable variation. [^9] A caveat on feature sets: the rule-gating model’s effective inputs include both the $12$-dimensional gate features $z(A)$ and the $24$-dimensional rule activity and recommendation indicators, which are nonlinear functions of the lottery primitives. The neural benchmarks operate on the raw $40$-dimensional encoding $\psi(A)$ with small, generic architectures. The comparison therefore does not hold features fixed; the feature-based baseline comparison in Appendix L (Table A.7) provides the closest feature-controlled evaluation.
Figure 1(a) visualizes the comparison with split-variability bands; rankings are stable across splits. Figure 1(b) reports a learning-curve diagnostic. Rule-gating outperforms MOT at small training fractions and remains essentially tied through moderate fractions; MOT pulls slightly ahead only above roughly $70\%$ of the training data, consistent with its larger number of free parameters requiring larger samples to pay off. Rule-gating consistently outperforms both the context-dependent and value-based networks across all fractions, confirming that a small rule repertoire captures variation beyond a single preference representation. Notably, rule-gating achieves strong performance even with limited data: at $10\%$ of the training sample it already outperforms the value-based network at full sample, reflecting the efficiency gains from the structured rule library.
The trial-weighted metric $\mathrm{MSE}_{w}$ (Table 3) preserves all pairwise orderings. Appendix Tables A.2 and A.3 report split-level comparisons and split-variability intervals.
Interpretability–accuracy tradeoff.
A gradient-boosted tree ensemble (GBT) trained on raw lottery summary statistics achieves MSE $0.0096$ (Appendix Table A.7), meaningfully below both rule-gating ($0.0117$) and MOT ($0.0114$). The interpretability cost of the RRM is therefore non-trivial. We view this as a worthwhile tradeoff because the RRM’s interpretability is structural, not post-hoc: for each menu, it decomposes the prediction into contributions from named behavioral procedures with economically interpretable weights. Post-hoc explainability tools identify which features matter but cannot decompose predictions into named procedures. This structural decomposition, together with the identification guarantees of Section 3, is what permits the mechanism-level analysis that follows.
Model Mean MSE SD Mean $\mathrm{MSE}_{w}$ Best LR
MOT 0.01139 0.00056 n.a. 0.001
Rule-gating 0.01168 0.00049 0.01155 0.010
Context-dependent 0.01344 0.00081 0.01330 0.010
Value-based 0.01921 0.00081 0.01904 0.010
Neural PT 0.02064 0.00085 0.02050 0.010
Neural CPT 0.02145 0.00093 0.02134 0.010
Neural EU 0.02215 0.00085 0.02205 0.010
Note: MOT learning rate is selected by test MSE following the published protocol of Peterson et al. (2021) (see Section 6.4); the HURD framework does not produce per-menu predictions, so $\mathrm{MSE}_{w}$ is unavailable for MOT in cross-validation.
(a) Mean test MSE across 50 random splits (split-variability band).
7.3 Responsibility weights $w_{f}$
Two patterns stand out. First, the attention rules A1 and A2 absorb the two largest shares of effective responsibility ($w_{\textsc{A1}}\approx 0.280$ and $w_{\textsc{A2}}\approx 0.260$), with a combined attention weight of about $54\%$. This is expected: the attention rules are decisive on essentially all menus and serve as a catch-all channel. Among the other rules, SAL (most salient comparison, $9.4\%$), REG (regret, $7.4\%$), MAP (modal payoff, $6.0\%$), MMx (aspiration, $5.2\%$), and MMn (security, $5.1\%$) carry the largest responsibility weights. Second, because multiple rules are typically active on the same menu (on average, about $5.8$ rules are both active and consistent with the majority choice direction), these responsibility weights reflect the outcome of competition among substitutable procedures rather than mechanical assignment to the only available rule.
It is useful to separate the library into two conceptual channels. The structured rules (MMn, MMa, MMx, MAP, SAL, SAL2, REG, REGmed, DIS, DISmed) encode distinct behavioral mechanisms that generate non-trivial, menu-dependent recommendations. The attention channel (A1 and A2) captures inattention, hasty responding, or trembling: A1 always recommends left and A2 always recommends right, regardless of lottery content. These attention rules absorb roughly half of effective responsibility, reflecting the prevalence of default-like behavior in the online (MTurk) environment. Because the marginal left-choice frequency is $\bar{p}\approx 0.510$ (essentially one-half), this weight cannot be explained by aggregate position bias; rather, the gate allocates attention weight in a menu-dependent way. Part of this weight may, however, reflect absorption of residual prediction error by always-active rules rather than a cleanly identified behavioral channel. Our interpretability claims concern the structured rules conditional on the inattention channel. Section 8.2 confirms that mechanism patterns persist when the attention channel is removed entirely.
The Herfindahl index $\mathrm{HHI}(F)=\sum_{f}w_{f}^{2}\approx 0.173$ yields an effective number of rules $N_{\mathrm{eff}}=1/\mathrm{HHI}\approx 5.8$, reflecting the concentration of responsibility on the attention channel; among the ten structured rules alone, weight is more evenly spread.
7.4 Rule concentration
(a) Predictive ablation index $\phi(f)$.
(b) Concentration Impact $\sigma_{N}(f)$.
We next quantify which procedures are necessary for predictive fit. Figure 3(a) reports the ablation index $\phi(f)$, the relative change in test MSE when rule $f$ is removed from the full library and the gate is retrained on the reduced set. We ablate each of the 10 behavioral rules (excluding attention rules A1/A2, which serve a structural role ensuring decisiveness at every menu). A small subset of rules accounts for most of the ablation impact: removing SAL2 or SAL produces the largest deterioration, followed by REGmed and REG. MAP, MMx, and MMn have moderate impacts, while DIS/DISmed have small impacts. Notably, SAL2 has a relatively high ablation impact despite carrying a lower average responsibility weight than SAL; this illustrates the distinction between $w_{f}$ (average usage) and $\phi(f)$ (marginal predictive contribution), and suggests that the second-most-salient comparison provides incremental content not replicated by other rules. MMa has a slightly negative ablation index, indicating that removing this low-weight rule marginally improves fit—a standard regularization effect when a near-redundant rule is dropped.
Ablations identify rules that matter for accuracy, but they do not indicate whether a rule acts as a unifier (concentrating effective responsibility on a small repertoire) or as a substitute that diversifies the fitted mixture. To capture this, in Figure 3(b), we report the relative change in the effective number of rules $\sigma_{N}(f)$. We find that $\sigma_{N}(f)$ is negative for all non-attention rules except REG, which has a small positive value. Removing REG slightly increases the effective number of rules, reflecting its specialized role: REG is active on only about $30\%$ of menus, and its removal frees weight to spread more evenly across the broadly active rules. For all other rules, $\sigma_{N}(f)<0$: removing a rule concentrates the mixture. The magnitude is largest for the same procedures that have large ablation impacts, indicating that these rules both attract mass and shape the overall mixture structure.
Overall, salience-based and regret-type comparisons account for the largest ablation impacts, as evidenced by their large $\phi(f)$, while the responsibility weights $w_{f}$ additionally highlight SAL and REG as the most prominent structured rules. Most remaining rules act as substitutes that diversify the fitted mixture, as indicated by their negative $\sigma_{N}(f)$ values. Appendix Table A.4 reports split-variability intervals for the absolute ablation effect $\Delta\mathrm{MSE}$, confirming that the main rankings are stable. More broadly, fold-level stability of the decomposition objects is reported in Appendix Table A.5: the rule rankings by $w_{f}$ are highly stable across independent training samples (mean Spearman $0.93$); $\phi(f)$ rankings show moderate stability (mean Spearman $0.44$), reflecting the greater sensitivity of ablation effects to sample composition. This decomposition provides a transparent characterization of which bounded-rationality mechanisms are empirically relevant in a large dataset of choice under risk—a level of structural interpretability not available from the more flexible predictors.
Library coverage
@@ 7.4 Rule concentration
Figure 4 shows that menus pinning down a unique decisive rule are rare ($\Pr[N_{\mathrm{cons}}(A)=1]\approx 0.16\%$), while the average number of rules both active and consistent with the majority direction is $\mathbb{E}[N_{\mathrm{cons}}(A)]\approx 5.8$, with about $9.5$ rules active per menu. The gate therefore learns selection among genuinely competing procedures.
@@
7.5 Comparative statics in rule responsibility
The figures in this section plot effective responsibility weights $\tilde{q}_{f}(A;\theta)$, which condition on rule activity (Equation (5)). Changes in $\tilde{q}_{f}$ across menus reflect two channels: (i) genuine shifts in the latent gate propensity $q_{f}(A;\theta)$ and (ii) mechanical denominator effects from the composition of the active set $\mathcal{F}_{A}$, which varies across menus. To verify that the patterns documented below are not artifacts of the activity normalization, Appendix Figures A.3A.4 reproduce the comparative statics using the latent gate outputs $q_{f}(A;\theta)$ directly (before conditioning on activity). The qualitative patterns are preserved: the latent gate itself shifts weight across rules in the same directions as the effective decomposition, confirming that the reported patterns reflect genuine variation in the learned gate rather than mechanical composition effects.
Complexity
and the associated excess dissimilarity
$d(A)\;:=\;\Delta_{\mathrm{CDF}}(A)-\big|EV(L^{1})-EV(L^{2})\big|.$
Figure 5 shows clear and interpretable shifts in rule selection along tradeoff complexity. As $\mathrm{TC}(A)$ increases, the two attention rules (A1/A2) display a pronounced increase in mean effective weight, consistent with the view that more complex menus are associated with a greater reliance on coarse inattention in the learned CA-RRM. Among the structured procedures, the most striking pattern is REG (regret): it carries its highest weight in the lowest-complexity deciles and drops steeply to near zero at high complexity, indicating that regret-type evaluation is concentrated on simple menus where state-by-state comparison is tractable. SAL (salience) rises with complexity, while MAP (modal payoff) displays an inverted-U pattern, peaking at moderate complexity and declining at both extremes. MMx (aspiration) shows a mild increase across deciles. Overall, increasing tradeoff complexity tilts the fitted mixture away from evaluative comparisons (regret, modal outcome) and toward attention-based procedures and salience.
Risk
A second natural driver is the risk profile of the menu. Different procedures emphasize different aspects of risk, such as dispersion, downside exposure, and sensitivity to rare events. As a first-pass risk descriptor available directly from the menu, we consider the absolute asymmetry in dispersion across the two lotteries,
$\mathrm{RiskAsym}(A)\;:=\;|\sigma(L^{1})-\sigma(L^{2})|,$
where $\sigma(L)$ denotes the standard deviation of lottery $L$. This measures how different the two lotteries are in terms of spread, regardless of which lottery is riskier. Figure 6 plots mean responsibility weights across deciles of this measure.
Figure 6 provides a complementary view based on dispersion asymmetry. The most striking pattern involves REG (regret): its weight is highest when the two lotteries have similar dispersion (low asymmetry) and declines steeply as the gap in spread widens. By contrast, the attention defaults A1 and A2 increase monotonically with risk asymmetry, suggesting that large differences in spread push the gate toward inattention. SAL and MMx show mild increases, while MAP declines gently. Thus, in menus where the two options differ markedly in spread, the gate shifts mass away from regret-type evaluation and toward attention defaults.
These diagnostics show that rule selection varies systematically with salient menu primitives: along complexity, the mixture tilts toward inattention and salience-based comparison and away from regret-type and modal-outcome evaluation; along dispersion asymmetry, regret-type comparison declines steeply while attention defaults rise. Menu-dependent selection across a small repertoire of procedures is economically meaningful.
7.6 Completeness and Restrictiveness
We supplement the absolute MSE comparisons with two diagnostics: completeness (how much of a flexible benchmark’s variation is captured by a structured model) and restrictiveness (how tightly a model class constrains the mappings it can fit). Together, they help separate performance gains reflecting substantive structure from gains due to sheer flexibility.
ML Completeness Score
Following Peysakhovich and Naecker (2017), Fudenberg and Puri (2022), and Fudenberg et al. (2022), we define the ML completeness of model $g$ relative to Neural EU (baseline) and MOT (flexible benchmark):
By construction, $\mathrm{Comp}(g_{\mathrm{EU}})=0$ and $\mathrm{Comp}(g_{\mathrm{MOT}})=1$; values in $(0,1)$ indicate partial recovery. Figure 7(a) shows that rule-gating recovers approximately $97\%$ of the baseline-to-MOT gap, compared to $\approx 27\%$ for the Value-based architecture, indicating that the identified decomposition captures nearly all of the structure captured by the most flexible benchmark.
Restrictiveness
(a) ML Completeness Scores $\mathrm{Comp}(g)$
(b) Restrictiveness $\mathrm{Restr}(g)$
Figure 7: Completeness and Restrictiveness.
Results.
Figure 7(b) shows that MOT is markedly less restrictive ($\mathrm{Restr}\approx 0.2$), while rule-gating and the neural utility benchmarks cluster near $1$. Figure 8 plots models in the (completeness, restrictiveness) plane: rule-gating occupies the upper-right region, capturing a large share of the variation captured by flexible benchmarks while retaining strong discipline against fitting arbitrary mappings.
7.7 Portability
To assess portability, we evaluate models trained on choices13k (Peterson et al., 2021) on the independent CPC18 dataset (Erev et al., 2017; Plonsky et al., 2017). [^10] The data can be downloaded at https://cpc-18.com/data/. We restrict to risk problems, yielding $686$ subjects, $|\mathcal{A}^{\mathrm{CPC}}|=180$ distinct menus, and 354,400354{,}400 total trials. [^11] In addition to menu-level aggregation, we also report a trial-weighted evaluation.
For each model, we freeze choices13k-estimated parameters and evaluate on CPC18 menus using
The outcome rescaling factor is computed from choices13k and applied identically to CPC18.121212Both datasets use a common monetary unit, so the rescaling simply puts outcomes on a common numerical scale. No hyperparameters are tuned on CPC18. Because CPC18 records individual-level trial data, we also evaluate at the trial level using the Brier score and Bernoulli log-loss.
Portability results.
Table 4 reports results. MOT—the best in-sample performer—deteriorates sharply under transfer, suggesting that much of its advantage is dataset-specific. Rule-gating retains strong predictive content: CPC18 MSE $0.0148$, ahead of the context-dependent model ($0.0173$) and far ahead of the structured neural benchmarks ($\approx 0.032$). Rule-gating therefore offers the best portability among all models with an interpretable behavioral decomposition.
Rule-gating 0.1952 0.5705 0.0148
Context-dependent 0.1977 0.5763 0.0173
Neural CPT 0.2111 0.6111 0.0319
Neural EU 0.2111 0.6115 0.0318
Neural PT 0.2120 0.6135 0.0325
Value-based 0.2122 0.6146 0.0323
MOT 0.2713 0.7513 0.0943
7.8 Two-step econometric estimates
Table 5 reports the results of the two-step econometric estimator described in Section 5, applied to the same data. The first stage groups menus into $47$ cells by $k$-means on $z(A)$ and recovers cellwise normalized weights $\hat{\omega}(x^{(k)})$ via constrained least squares on the $H^{(k)}$ matrices (minimizing $\|H^{(k)}\omega\|^{2}$ subject to $\omega_{f_{0}}=1$, $\omega_{f}\geq\varepsilon$). The second stage regresses $\log\hat{\omega}_{f}(x^{(k)})$ on the cell centroids $1,x^{(k)}]$ to recover the affine gate coefficients $(\hat{\alpha}_{f},\hat{\beta}_{f})$. This implementation involves two practical modifications relative to the estimator analyzed in Theorem 3: approximate binning via $k$-means in place of exact common-feature cells, and positivity constraints on the first-stage weights. Both are standard: binning is asymptotically innocuous under smooth variation of $\omega(x)$, and the positivity constraints do not bind when true weights are strictly positive. The two-step results should accordingly be read as an econometric cross-check motivated by the identification theory, rather than as carrying the full inferential status of Theorem 3.
The two-step responsibility weights $w_{f}$ show moderate-to-good agreement with the MSE-fitted decomposition: Spearman rank correlation is $0.70$ ($\text{Pearson}=0.40$). SAL, REG, and A2 appear in the top five under both estimators, but point estimates and ordinal ranks differ. The most notable discrepancy involves the attention rules: A1 carries the largest MSE weight ($w_{f}=0.28$) but drops to fifth under the two-step procedure ($w_{f}=0.11$), while SAL is the most prominent two-step rule ($w_{f}=0.23$) but ranks third under MSE ($w_{f}=0.09$). These differences are expected. The attention rules are one-sided: A1 always recommends $L^{1}$ and A2 always recommends $L^{2}$, regardless of the menu. They therefore contribute no within-cell variation in decisive-side patterns — precisely the variation that the first-stage constrained least squares exploits to identify rule weights. Their two-step weights are identified only residually, after the structured rules with genuine recommendation variation have been fit. By contrast, the MSE estimator minimizes prediction error globally and can freely load weight onto the inattention rules.
The overidentification $J$-test (Table 5, last column) rejects the affine gate restriction at the $1\%$ level for most rules, indicating that the linear-index parameterization $\log\omega_{f}(x)=\alpha_{f}+\beta_{f}^{\top}z$ is too parsimonious to capture the full pattern of cellwise weight variation. A1 ($J=26.1$, $p=0.86$) and MMa ($J=45.4$, $p=0.11$) are the exceptions: the former is consistent with attention being approximately constant across feature values, while the latter carries negligible weight under both estimators. This rejection is not unexpected: with 47 cells and $d_{\mathrm{eff}}+1=12$ parameters, the test has 35 degrees of freedom and substantial power against smooth departures from linearity.
Both procedures share the same functional form: $\log(q_{f}(x)/q_{f_{0}}(x))=\alpha_{f}+\beta_{f}^{\top}x$; the softmax is the normalized-probability representation of the affine index, not a separate gate class. The distinction is in the estimator, not the model: the two-step recovers cellwise weights nonparametrically and then fits the affine index via WLS, exposing the overidentifying restrictions that the $J$-test evaluates, while the MSE procedure optimizes prediction directly over the same parameterization without testing those restrictions. When the $J$-test rejects—as it does for most rules—the two estimators diverge in interpretation: the two-step estimator recovers cellwise weights with a specification diagnostic, while the MSE estimator yields a pseudo-true predictive projection. The MSE decomposition remains informative as a descriptive summary but should not be read as the exactly identified structural gate unless the affine restriction is supported. Importantly, Theorem 1 is fully nonparametric and does not require an affine gate: the cellwise identification of relative weights holds for any functional form. The $J$-test rejection therefore suggests that a richer second-step specification—such as a sieve or kernel parameterization—could reduce or eliminate rejection while preserving the first-step identification logic. This is a natural direction for future work.
Table 5: Two-step econometric estimator vs. MSE-fitted responsibility weights $w_{f}$. The two-step estimator recovers cellwise rule weights by constrained least squares on $H^{(k)}$, then maps log-weights into affine gate parameters. Bootstrap standard errors (100 resamples) in parentheses. The $J$-statistic tests the affine gate restriction ($\chi^{2}_{35}$).
Rule $w_{f}$ (two-step) $w_{f}$ (MSE) Difference $J$-stat
MMn 0.131 (0.062) 0.051 +0.080+0.080 293.2∗∗∗
MMa 0.000 (0.012) 0.008 0.008-0.008 45.4
MMx 0.170 (0.063) 0.052 +0.118+0.118 252.0∗∗∗
MAP 0.023 (0.062) 0.060 0.037-0.037 357.6∗∗∗
SAL 0.228 (0.070) 0.094 +0.134+0.134 197.0∗∗∗
SAL2 0.072 (0.034) 0.039 +0.033+0.033 210.1∗∗∗
REG 0.099 (0.030) 0.074 +0.025+0.025 152.2∗∗∗
REGmed 0.039 (0.051) 0.032 +0.007+0.007 212.3∗∗∗
DIS 0.004 (0.022) 0.017 0.013-0.013 186.5∗∗∗
DISmed 0.001 (0.028) 0.033 0.032-0.032 136.4∗∗∗
A1 0.109 (0.070) 0.280 0.171-0.171 26.1
A2 0.123 (0.043) 0.260 0.137-0.137
Pearson correlation: 0.401; Spearman rank correlation: 0.699
$^{***}p<0.01$; $H_{0}$: $\log\omega_{f}(x)$ is affine in $x$
8 Further Robustness
We subject the framework to targeted robustness checks addressing library composition, the attention channel, the activity discipline, and the choice of gate features.